# EvidenceMap: Learning Evidence Analysis to Unleash the Power of Small Language Models for Biomedical Question Answering

### v2

@@ 6.2 Textual Knowledge Augmented LLM
Textual evidence, as an external knowledge, can significantly enhance the quality of language model generation in professional domains. Methods for knowledge augmentation can be categorized into two types. The first is the augmentation during inference. RAG Lewis et al. (2020) and Self-RAG Asai et al. (2023) belong to this category, which improve generation effectiveness by retrieving relevant documents as factual knowledge and adjusting these texts during inference. Meanwhile, a model can reduce hallucinations from generation by incorporating reasoning knowledge, as demonstrated by CoT Wei et al. (2022) and ToT Yao et al. (2024), a series of sequential or structured reasoning steps are generated and executed explicitly. Furthermore, methods with an augmentation mixture of factual and reasoning knowledge can further combine the strengths of both, such as RAT Wang et al. (2024). Another type of augmentation is by tuning the model with domain text. Soft prompt tuning Lester et al. (2021); Yang et al. (2023) uses implicit learnable vectors to replace explicit tokens as prompts to flexibly adapt to multiple downstream tasks. Instruction tuning Wang et al. (2023b); Lin et al. (2023) improves models in knowledge-intensive tasks by constructing specific instruction datasets and learning the utilization of input information and contextual understanding. Compared to the aforementioned methods, our study proposes a novel knowledge augmentation process that prompts the inference of generative models by training the explicit evidence analysis abilities with a small pre-trained language model.
@@ 7 Conclusion
In this work, we present a novel framework for generative biomedical question answering named EvidenceMap, which efficiently addresses the robustness issues of language model generation with multiple pieces of evidence. Inspired by evidence analysis in the biomedical domain, our framework explicitly trains a small language model on multifaceted aspects of evidence to prompt the generation of answers. The experimental results and representative cases on public datasets show that EvidenceMap can significantly surpass the RAG method with equivalent and larger language models by fine-tuning a 66M pre-trained model to prompt a generative model around 3B.
@@
Third, despite our method demonstrating superior performance in small language models, we will continue to conduct in-depth research on the interpretability and visualization of model inference, exploring further advantages in evidence analysis.
Fourth, the generalizability of learning evidence analysis and the potential transferability to other models still remains an area for further exploration.
@@ Appendix E Results of Language Fluency
We explore the fluency of language generated by our framework. As stated in Appendix A.3, an LLM is employed to score accuracy and fluency of the generated output at the same time. The average values of all samples from the test set are calculated and are shown in Table 6. The results indicate that our framework achieves the best and competitive fluency quality on BioASQ and PubMedQA, respectively.

### v4

@@ 6.2 Textual Knowledge Augmented LLM
Textual evidence, as an external knowledge, can significantly enhance the quality of language model generation in professional domains. Methods for knowledge augmentation can be categorized into two types. The first is the augmentation during inference. RAG Lewis et al. (2020) and Self-RAG Asai et al. (2023) belong to this category, which improve generation effectiveness by retrieving relevant documents as factual knowledge and adjusting these texts during inference. Meanwhile, a model can reduce hallucinations from generation by incorporating reasoning knowledge, as demonstrated by CoT Wei et al. (2022) and ToT Yao et al. (2024), a series of sequential or structured reasoning steps are generated and executed explicitly. Furthermore, methods with an augmentation mixture of factual and reasoning knowledge can further combine the strengths of both, such as RAT Wang et al. (2024). Another type of augmentation is by tuning the model with domain text. Soft prompt tuning Lester et al. (2021); Yang et al. (2023) uses implicit learnable vectors to replace explicit tokens as prompts to flexibly adapt to multiple downstream tasks. Instruction tuning Wang et al. (2023); Lin et al. (2023) improves models in knowledge-intensive tasks by constructing specific instruction datasets and learning the utilization of input information and contextual understanding. Compared to the aforementioned methods, our study proposes a novel knowledge augmentation process that prompts the inference of generative models by training the explicit evidence analysis abilities with a small pre-trained language model.
@@ 7 Conclusion
In this work, we present a novel framework for generative biomedical question answering named EvidenceMap, which efficiently addresses the robustness issues of language model generation with multiple pieces of evidence. Inspired by evidence analysis in the biomedical domain, our framework fine-tunes a small pre-trained language model to learn the analysis of multifaceted aspects of evidence and prompt a small generative model to produce answers. The experimental results and representative cases on public datasets show that EvidenceMap can significantly surpass popular reasoning methods with equivalent and larger language models by fine-tuning a pre-trained model with only 66M parameters.
@@
Third, despite our method demonstrating superior performance in generation with small language models, we will continue to conduct in-depth research on the interpretability and stability of model inference, exploring further advantages in evidence analysis, and with larger language models.
Fourth, the generalizability of learning evidence analysis and the potential transferability to various pre-trained models still remains an area for further exploration, which will justify whether evidence analysis is a universal competency, and further improve the efficiency of applying evidence analysis.
@@ Appendix E Results of Language Fluency
We explore the fluency of the language generated by our framework to verify whether the method preserves the original language generation capabilities. As stated in Appendix A.3, an LLM is employed to score accuracy and fluency of the generated output at the same time. The average values of all samples from the test set are calculated and are shown in Table 6. The results indicate that our framework achieves the highest and competitive fluency on BioASQ and PubMedQA, respectively, while improving the quality of the answers.
@@
Appendix F Complexity Analysis
The complexity of the framework arises primarily from the combinatorial analysis of multiple evidence during training. Specifically, if there are $N$ pieces of evidence, it is necessary to evaluate the supportive relationship of each piece of evidence to the question, which involves representing $N$ relationships. Furthermore, it is essential to determine the logical relationships between two pieces of evidence, resulting in the representation of $\frac{N\times(N-1)}{2}$ relationships. In practice, we set the number of evidence as 5 and the batch size as 4, which allows the entire training process to run on a single GPU. In addition, conducting evidence quality assessment sequentially or selecting more efficient parallel encoders can further reduce complexity.